# Singularities of solutions of nonlocal nonlinear equations

### v1

@@ 3.2. Interior regularity estimates
for some $C=C(n,s,p,\Lambda,\varepsilon,(\mathrm{diam}\,\Omega)^{\varepsilon}\|b_{1}\|_{L^{n/(sp-\varepsilon)}(\Omega)})>0$, where $\mathrm{DG}^{s,p}$ is the fractional De Giorgi class given in [15, Section 6.1]. Thus, the desired result follows from [15, Theorem 6.4]. ∎
@@
Since (3.12) holds with $\Omega=B_{R}$, the desired result follows from [15, Theorem 6.9]. ∎
@@
Let $u$ be a weak solution of (1.1) in $B_{R}\setminus\{0\}$ such that $u\geq 0$ in $B_{R}\setminus\{0\}$. Then for any $r\leq R/2$,
@@
If we let $j\in\{1,\dots,N\}$ be such that
$\inf_{B_{r}\setminus B_{r/2}}{u}=\inf_{B_{r/4}(y_{j})}{u},$
then
@@ 4. Removable singularity
We first prove that $u\in W^{s,p}(B_{R/2})$. Since $E$ has $(t,q)$-capacity zero, Lemma 2.17 shows that there exists a sequence of smooth functions $\eta_{j}$ such that $0\leq\eta_{j}\leq 1$ in $\mathbb{R}^{n}$, $\eta_{j}=1$ in a neighborhood of $E$ and $[\eta_{j}]_{W^{t,q}(\mathbb{R}^{n})}\to 0$ as $j\to\infty$. We may assume that $\eta_{j}\to 0$ a.e. in $\mathbb{R}^{n}$. Let $\bar{\eta}_{j}=1-\eta_{j}$. Then $\bar{\eta}_{j}$ is a nonnegative smooth function vanishing in a neighborhood of $E$ such that $[\bar{\eta}_{j}]_{W^{t,q}(\mathbb{R}^{n})}\to 0$ and $\bar{\eta}_{j}\to 1$ a.e. in $\mathbb{R}^{n}$ as $j\to\infty$. Let $R/2\leq\rho<r\leq R$, $\beta_{0}=\delta(p-1)>0$ and
@@
Note that $|E|=0$ by Lemma 2.15. Thus, Lemma 3.3 with $\bar{\eta}=\bar{\eta}_{j}$ shows that
@@ 5. Isolated singularity
The goal of this section is to describe the asymptotic behavior of a singular solution near an isolated non-removable singularity. Namely, we provide the proofs of Theorem 1.3 and Theorem 1.5. In the rest of the paper, we always assume that $\Omega$ contains the origin and that $\Omega$ is bounded as Theorem 1.3 and Theorem 1.5 focus on the isolated singularity at the origin. Moreover, we may assume that $u\in C(\overline{\Omega}\setminus\{0\})\cap W^{s,p}(\Omega\setminus B_{\rho})$ for all $\rho>0$ by recalling Theorem 3.7 and choosing $\Omega^{\prime}\Subset\Omega$ if necessary. The last assumption allows us to avoid technical issues arising from the regularity of $u$ near $\partial\Omega$.
@@
Since $\{0\}$ is of $(s,p)$-capacity zero, there exists a sequence of smooth functions $\bar{\eta}_{j}$ such that $0\leq\bar{\eta}_{j}\leq 1$ in $\mathbb{R}^{n}$, $\lim_{j\to\infty}[\bar{\eta}_{j}]_{W^{s,p}(\mathbb{R}^{n})}=0$, $\bar{\eta}_{j}\to 1$ a.e. in $\mathbb{R}^{n}$ as $j\to\infty$ and $\bar{\eta}_{j}$ vanishes in a neighborhood of $\{0\}$; see the beginning of the proof of Theorem 1.1. We now set $E=\{0\}$, let $B_{r_{1}}(x_{0})\Subset B_{r_{2}}(x_{0})\subset B_{R}$ and take $\bar{\eta}=\bar{\eta}_{j}$ in (3.11). Since $w_{-}\coloneqq(u-k)_{-}\in L^{\infty}(B_{R})$ for all $k\in\mathbb{R}$ by the assumption, it follows from Fatou’s lemma that
@@
for some $C=C(n,s,p,\Lambda,\varepsilon,R^{\varepsilon}\|b_{1}\|_{L^{n/(sp-\varepsilon)}(B_{R})})>0$, where $\mathrm{DG}_{-}^{s,p}$ is the fractional De Giorgi class given in [15, Section 6.1]. Thus, the desired estimate follows from [15, Proposition 6.8]. ∎
@@
Let $B_{R}=B_{R}(0)\subset\Omega$. For any $M>0$, there exists a small radius $r=r_{M}\in(0,R/2)$ such that
$u\geq M\quad\text{in $B_{2r}\setminus B_{r/2}$},$
by the Harnack inequality (Theorem 3.9) and (5.1). By using Lemma 5.2 and the fact that $u\geq 0$ in $B_{R}\setminus\{0\}$, we obtain that
@@
for some $p_{0}\in(0,1)$. By combining all estimates above, we conclude that $M\leq CL+C$ for some $C>0$, which leads to the contradiction. ∎
In order to prove the upper bounds in Theorem 1.3 and Theorem 1.5, we first introduce the following invariant.
@@
where $C$ is a positive constant depending on $n$, $s$, $p$, $\Lambda$, $R$, $d$, $\eta$ and $u$. Since the function $\varphi$ is admissible for $\mathrm{cap}_{s,p}(\overline{B_{r}}\setminus B_{r/2},\Omega)$, it follows from Lemma 2.13 that
(5.16) $\Lambda\mathcal{E}(\varphi,\varphi)\geq[\varphi]_{W^{s,p}(\mathbb{R}^{n})}^{p}\geq\mathrm{cap}_{s,p}(\overline{B_{r}}\setminus B_{r/2},\Omega)\simeq\begin{cases}r^{n-sp}&\text{if}~{}sp<n,\\ (-\log r)^{1-p}&\text{if}~{}sp=n.\end{cases}$
Thus, by taking $r\in(0,1)$ sufficiantly small, we obtain from (5.15) and (5.16) that
@@
Therefore, the desired upper bound (5.5) for sufficiently small $r>0$ follows from (5.17) and the Harnack inequality (Theorem 3.9). ∎
@@
Therefore, the conclusion follows from the same argument provided in the proof of [34, Lemma 2]. ∎
The upper bound of singular solutions is now an immediate consequence of Lemma 5.4 and Lemma 5.5, both in the subcritical and critical cases.
Proofs of the upper bound in Theorem 1.3 and Theorem 1.5.
We may assume that $u$ is positive in $\Omega\setminus\{0\}$ by adding a suitable constant $C$ to $u$. In this case, $v\coloneqq u+C$ satisfies $\mathcal{L}v+\bar{b}(x,v)=0$ in $\Omega\setminus\{0\}$, where $\bar{b}(x,z)\coloneqq b(x,z-C)$ enjoys the same structure condition as $b$, namely,
$\bar{b}(x,z)\leq\max\{2^{p-2},1\}b_{1}(x)|z|^{p-1}+C^{p-1}\max\{2^{p-2},1\}b_{1}(x)+b_{2}(x)\eqqcolon\bar{b}_{1}(x)|z|^{p-1}+\bar{b}_{2}(x).$
Therefore, we can apply Lemma 5.4 and Lemma 5.5, which leads us to the desired upper bound of singular solutions. ∎
@@
where $m=m_{r}$ is given as in (5.4). Once we prove the claim (5.20), then it follows from the Harnack inequality (Theorem 3.9) that
$\sup_{B_{r}\setminus B_{r/2}}{u}\leq Cr^{-\tau(1-\delta)}+C\left(\frac{r}{R}\right)^{\frac{sp}{p-1}}\mathrm{Tail}(u_{-};0,R)+C\left(\frac{r}{R}\right)^{\frac{\varepsilon}{p-1}}\left(R^{\varepsilon}\|b_{2}\|_{L^{n/(sp-\varepsilon)}(B_{R})}\right)^{\frac{1}{p-1}}.$
Hence, Corollary 1.2 shows that $u$ has a removable singularity at the origin, yielding a contradiction.
@@
By recalling (5.16), we have
@@
Let $\phi$ be the $(-\Delta_{p})^{s}$-potential of $\overline{B_{r}}$ in $B_{\delta R}$, as defined in Kim–Lee–Lee [22, Definition 2.15], then $\phi$ satisfies the conditions (for $\varphi$) in Lemma 5.6. Thus, Lemma 5.6 and Lemma 5.7 show that the value $K$ defined by (5.18) is a positive constant. We write
@@
$\left|\int_{\Omega}b(x,u(x))\phi(x)\,\mathrm{d}x\right|\leq\int_{B_{\delta R}}|b(x,u(x))|\,\mathrm{d}x\leq C(\delta R)^{\varepsilon}\leq\frac{K}{8}$
@@
$[\phi]_{W^{s,p}(\mathbb{R}^{n})}^{p}=\mathrm{cap}_{s,p}(\overline{B_{r}},B_{\delta R})\simeq\left(r^{-\frac{n-sp}{p-1}}-R^{-\frac{n-sp}{p-1}}\right)^{1-p}\lesssim r^{n-sp},$
@@
(5.23) $\widetilde{M}=\max_{\overline{B_{R}}\setminus B_{\delta r}}u\gtrsim r^{-(n-sp)/(p-1)}.$
It only remains to transport this information to $u$ on $\partial B_{r}$. For this purpose, choose $\rho\in[\delta r,R]$ so that
@@
and so the Harnack inequality (Theorem 3.9) finishes the proof. ∎
@@
As in the proof of the upper bound, we may assume that $u$ is positive in $\Omega\setminus\{0\}$. Thus, the desired follows from Lemma 5.8. ∎

### v2

@@ 3.2. Interior regularity estimates
for some $C=C(n,s,p,\Lambda,\varepsilon,(\mathrm{diam}\,\Omega)^{\varepsilon}\|b_{1}\|_{L^{n/(sp-\varepsilon)}(\Omega)})>0$, where $\mathrm{DG}^{s,p}$ is the fractional De Giorgi class given in [17, Section 6.1]. Thus, the desired result follows from [17, Theorem 6.4]. ∎
@@
Since (3.12) holds with $\Omega=B_{R}$, the desired result follows from [17, Theorem 6.9]. ∎
@@
Assume $n\geq 2$. Let $u$ be a weak solution of (1.1) in $B_{R}\setminus\{0\}$ such that $u\geq 0$ in $B_{R}\setminus\{0\}$. Then for any $r\leq R/2$,
@@
Let $j\in\{1,\dots,N\}$ be such that
$\inf_{B_{r}\setminus B_{r/2}}{u}\geq\inf_{B_{r/4}(y_{j})}{u}.$
Then, since $B_{r}\setminus B_{r/2}$ is connected, we obtain that
@@ 4. Removable singularity
We first prove that $u\in W^{s,p}(B_{R/2})$. Since $E$ has $(t,q)$-capacity zero, so is $E\cap\overline{B}_{R}$. Since $E\cap\overline{B}_{R}$ is compact, by (2.8), for each $j\in\mathbb{N}$, there are open sets $G_{j}$ and $G_{j}^{\prime}$ such that
$E\cap\overline{B}_{R}\Subset G_{j}\Subset G_{j}^{\prime}\Subset B_{2R}\quad\text{and}\quad\mathrm{cap}_{t,q}(G_{j}^{\prime},B_{2R})<1/j.$
Then $\mathrm{cap}_{t,q}(\overline{G}_{j},B_{2R})<1/j$, and hence by definition there are $\eta_{j}^{\prime}\in C^{\infty}_{c}(B_{2R})$ such that $\eta_{j}^{\prime}\geq 1$ on $\overline{G}_{j}$ and $[\eta_{j}^{\prime}]_{W^{t,q}(\mathbb{R}^{n})}<1/j$. Thus, there is a mollification $\eta_{j}\in C^{\infty}_{c}(B_{2R})$ of $\min\{\eta_{j}^{\prime},1\}_{+}$ such that $0\leq\eta_{j}\leq 1$ in $\mathbb{R}^{n}$, $\eta_{j}=1$ in a neighborhood of $E\cap\overline{B}_{R}$ and $[\eta_{j}]_{W^{t,q}(\mathbb{R}^{n})}<1/j$. We may also assume that $\eta_{j}\to 0$ a.e. in $\mathbb{R}^{n}$ after taking a subsequence if necessary. Let $\bar{\eta}_{j}=1-\eta_{j}$. Then $\bar{\eta}_{j}$ is a nonnegative smooth function vanishing in a neighborhood of $E\cap\overline{B}_{R}$ such that $[\bar{\eta}_{j}]_{W^{t,q}(\mathbb{R}^{n})}\to 0$ and $\bar{\eta}_{j}\to 1$ a.e. in $\mathbb{R}^{n}$ as $j\to\infty$.
Let $R/2\leq\rho<r\leq R$, $\beta_{0}=\delta(p-1)>0$ and
@@
Note that $|E\cap\overline{B}_{R}|=0$ by Lemma 2.14. Thus, Lemma 3.3 with $E$ and $\bar{\eta}$ replaced by $E\cap\overline{B}_{R}$ and $\bar{\eta}_{j}$, respectively, shows that
@@ 5. Isolated singularity
The goal of this section is to describe the asymptotic behavior of a singular solution near an isolated non-removable singularity. Namely, we provide the proofs of Theorems 1.3 and 1.5. In the rest of the paper, we always assume that $\Omega$ contains the origin and that $\Omega$ is bounded as Theorems 1.3 and 1.5 focus on the isolated singularity at the origin. Note, however, that these theorems still hold for unbounded $\Omega$. Moreover, we may assume that $u\in C(\overline{\Omega}\setminus\{0\})\cap W^{s,p}(\Omega\setminus B_{\rho})$ for all $\rho>0$ by recalling Theorem 3.7 and choosing $\Omega^{\prime}\Subset\Omega$ if necessary. The last assumption allows us to avoid technical issues arising from the regularity of $u$ near $\partial\Omega$.
@@
Since $\{0\}$ is of $(s,p)$-capacity zero, there exists a sequence of smooth functions $\bar{\eta}_{j}$ such that $0\leq\bar{\eta}_{j}\leq 1$ in $\mathbb{R}^{n}$, $\lim_{j\to\infty}[\bar{\eta}_{j}]_{W^{s,p}(\mathbb{R}^{n})}=0$, $\bar{\eta}_{j}\to 1$ a.e. in $\mathbb{R}^{n}$ as $j\to\infty$ and $\bar{\eta}_{j}$ vanishes in a neighborhood of $\{0\}$; see the beginning of the proof of Theorem 1.1. We now set $E=\{0\}$, let $B_{r_{1}}(x_{0})\Subset B_{r_{2}}(x_{0})\subset B_{R}$ and take $\bar{\eta}=\bar{\eta}_{j}$ in (3.11). Since $w_{-}\coloneqq(u-k)_{-}\in L^{\infty}(B_{R})$ for all $k\in\mathbb{R}$ by assumption, it follows from Fatou’s lemma that
@@
for some $C=C(n,s,p,\Lambda,\varepsilon,R^{\varepsilon}\|b_{1}\|_{L^{n/(sp-\varepsilon)}(B_{R})})>0$, where $\mathrm{DG}_{-}^{s,p}$ is the fractional De Giorgi class given in [17, Section 6.1]. Thus, the desired estimate follows from [17, Proposition 6.8]. ∎
@@
Let $B_{R}=B_{R}(0)\subset\Omega$. Then, for each $M>0$, there exists a radius $r=r_{M}\in(0,R/2)$ such that
$u\geq M$
in $B_{r}\setminus B_{r/2}$ for $n\geq 2$ and in $(r/2,r)$ or in $(-r,-r/2)$ for $n=1$, by the Harnack inequality (Theorem 3.9 for $n\geq 2$ and Theorem 3.8 for $n=1$) and (5.1). By using Lemma 5.2 and the fact that $u\geq 0$ in $B_{R}\setminus\{0\}$, we obtain that
@@
for some $p_{0}\in(0,1)$. By combining all the estimates above and taking sufficiently small $r$, we conclude that $M\leq CL+C$ for some $C>0$, which leads to the contradiction. ∎
In order to prove the upper bounds in Theorems 1.3 and 1.5, we first introduce the following invariant.
@@
where $C$ is a positive constant depending on $n$, $s$, $p$, $\Lambda$, $R$, $d$, $\eta$ and $u$. Since the function $\varphi$ is admissible for $\mathrm{cap}_{s,p}(\overline{B}_{r}\setminus B_{r/2},B_{3R})$, it follows from Lemma 2.12 that
(5.16) $\Lambda\mathcal{E}(\varphi,\varphi)\geq[\varphi]_{W^{s,p}(\mathbb{R}^{n})}^{p}\geq\mathrm{cap}_{s,p}(\overline{B}_{r}\setminus B_{r/2},B_{3R})\eqsim\begin{cases}r^{n-sp}&\text{if}~{}sp<n,\\ (-\log r)^{1-p}&\text{if}~{}sp=n,\end{cases}$
provided that $r$ is sufficiently small. Thus, by taking small $r\in(0,1)$, we obtain from (5.15) and (5.16) that
@@
Therefore, the desired upper bound (5.5) for sufficiently small $r>0$ follows from (5.17) and the Harnack inequality (Theorem 3.9).
Next, we assume that $n=1$. In this case, the argument proceeds analogously to the case $n\geq 2$, with $M$ and $m$ in (5.4) now defined by
(5.18) $M=M_{r}^{\pm}=\max_{I_{r}^{\pm}}u\quad\text{and}\quad m=m_{r}^{\pm}=\min_{I_{r}^{\pm}}u,$
where
(5.19) $I_{r}^{+}=[r/2,r]\quad\text{and}\quad I_{r}^{-}=[-r,-r/2].$
The only difference is that the function $\varphi$ is now admissible for $\mathrm{cap}_{s,p}(I_{r}^{\pm},(-3R,3R))$ instead of $\mathrm{cap}_{s,p}(\overline{B}_{r}\setminus B_{r/2},B_{3R})$. Nevertheless, by Lemma 2.12, we still obtain the lower bound
(5.20) $\mathrm{cap}_{s,p}(I_{r}^{\pm},(-3R,3R))\gtrsim\begin{cases}r^{n-sp}&\text{if}~{}sp<n,\\ (-\log r)^{1-p}&\text{if}~{}sp=n,\end{cases}$
which replaces (5.16). This leads to (5.17) with the new $m=m_{r}^{\pm}$, and consequently (5.5) with the new $M=M_{r}^{\pm}$ also follows from the Harnack inequality in Theorem 3.8. ∎
@@
Therefore, the conclusion follows from the same argument provided in the proof of [40, Lemma 2]. ∎
The upper bound of singular solutions is now an immediate consequence of Lemmas 5.4 and 5.5, both in the subcritical and critical cases.
Proofs of the upper bound in Theorems 1.3 and 1.5.
We may assume, by adding a suitable constant $C$ to $u$, that $u$ is positive in $\Omega\setminus\{0\}$. In this case, $v\coloneqq u+C$ satisfies $\mathcal{L}v+\bar{b}(x,v)=0$ in $\Omega\setminus\{0\}$, where $\bar{b}(x,z)\coloneqq b(x,z-C)$ enjoys the same structure condition as $b$, namely,
$\bar{b}(x,z)$ $\leq\max\{2^{p-2},1\}b_{1}(x)|z|^{p-1}+C^{p-1}\max\{2^{p-2},1\}b_{1}(x)+b_{2}(x)$
$\eqqcolon\bar{b}_{1}(x)|z|^{p-1}+\bar{b}_{2}(x).$
Therefore, we can apply Lemmas 5.4 and 5.5, which leads us to the desired upper bound of singular solutions. ∎
@@
where $m=m_{r}$ is given as in (5.4) if $n\geq 2$, and $m=m_{r}^{\pm}$ as in (5.18) if $n=1$. Once we prove the claim (5.23), then it follows from Theorem 3.9 that
$\sup_{B_{r}\setminus B_{r/2}}{u}\leq Cr^{-\tau(1-\delta)}+C\left(\frac{r}{R}\right)^{\frac{sp}{p-1}}\mathrm{Tail}(u_{-};0,R)+C\left(\frac{r}{R}\right)^{\frac{\varepsilon}{p-1}}\left(R^{\varepsilon}\|b_{2}\|_{L^{n/(sp-\varepsilon)}(B_{R})}\right)^{\frac{1}{p-1}}$
if $n\geq 2$, and from Theorem 3.8 and the nonnegativity assumption on $u$ that
$\sup_{I_{r}^{\pm}}{u}$ $\leq Cr^{-\tau(1-\delta)}+C\mathrm{Tail}(u_{-};3r/4,r/4)+C\left(r^{\varepsilon}\|b_{2}\|_{L^{n/(sp-\varepsilon)}(r/4,5r/4)}\right)^{\frac{1}{p-1}}$
$\leq Cr^{-\tau(1-\delta)}+C\left(\frac{r}{R}\right)^{\frac{sp}{p-1}}\mathrm{Tail}(u_{-};0,R)+C\left(\frac{r}{R}\right)^{\frac{\varepsilon}{p-1}}\left(R^{\varepsilon}\|b_{2}\|_{L^{n/(sp-\varepsilon)}(0,2R)}\right)^{\frac{1}{p-1}}$
if $n=1$, where $I_{r}^{\pm}$ is given in (5.19). Hence, Corollary 1.2 shows that $u$ has a removable singularity at the origin, yielding a contradiction.
@@
By recalling (5.16) for $n\geq 2$ and (5.20) for $n=1$, and applying Lemma 2 in Serrin [39], we have
@@
Let $\phi$ be the $(-\Delta_{p})^{s}$-potential of $\overline{B}_{r}$ in $B_{\delta R}$, as defined in Kim–Lee–Lee [26, Definition 2.15], then $\phi$ satisfies the conditions (for $\varphi$) in Lemma 5.6. Thus, Lemmas 5.6 and 5.7 show that the value $K$ defined by (5.21) is a positive constant. We write
@@
$|J|\leq\int_{B_{\delta R}}|b(x,u(x))|\,\mathrm{d}x\leq C(\delta R)^{\varepsilon}\leq\frac{K}{8}$
@@
$[\phi]_{W^{s,p}(\mathbb{R}^{n})}^{p}=\mathrm{cap}_{s,p}(\overline{B}_{r},B_{\delta R})\eqsim\left(r^{-\frac{n-sp}{p-1}}-R^{-\frac{n-sp}{p-1}}\right)^{1-p}\lesssim r^{n-sp},$
@@
(5.26) $\widetilde{M}\gtrsim r^{-(n-sp)/(p-1)}.$
It remains to transfer this information to the behavior of $u$ in an annulus whose inner and outer radii are comparable to $r$. We first assume that $n\geq 2$. We choose $\rho\in[\delta r,R]$ so that
@@
and so the Harnack inequality (Theorem 3.9) finishes the proof.
Next, we assume that $n=1$. In this case, we have
$\widetilde{M}=\max_{[\delta r,R]}u\quad\text{or}\quad\widetilde{M}=\max_{[-R,-\delta r]}u.$
Without loss of generality, we may assume that $\widetilde{M}=\max_{[\delta r,R]}u$. Then, by repeating the proof for the case $n\geq 2$ and using the Harnack inequality in Theorem 3.8 instead of Theorem 3.9, we conclude that
$u\gtrsim|x|^{-(1-sp)/(p-1)}$
in $[\delta r,\bar{c}r]$ for some $\bar{c}>0$. Moreover, the weak Harnack inequality developed in Lemma 5.2 shows that
$\left(\fint_{B_{\bar{c}r}}u^{p_{0}}\,\mathrm{d}x\right)^{\frac{1}{p_{0}}}\leq C\inf_{B_{\bar{c}r}}u+C\mathrm{Tail}(u_{-};0,R)+C\left(R^{\varepsilon}\|b_{2}\|_{L^{\frac{n}{sp-\varepsilon}}(B_{R})}\right)^{\frac{1}{p-1}}.$
The lower bound in the positive interval $[\delta r,\bar{c}r]$ is then transferred to the desired bound in the negative interval $[-\bar{c}r,-\delta r]$. ∎
@@
As in the proof of the upper bound, we may assume that $u$ is positive in $\Omega\setminus\{0\}$. Thus, the desired result follows from Lemma 5.8. ∎